OpenStudy (anonymous):

solve for x: log (7-x) - log (3x+2)=1. please help, have to take precal exam in the morning

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

subtraction=division

OpenStudy (anonymous):

that doesnt give me enough info bro haha im sloww. can you tell me the answer so i can go through and work out the problem?

OpenStudy (anonymous):

The answer, if I knew it, would help you less. log equations require steps. the left hand side can be grouped in to one term log a - log b= log (a/b)

OpenStudy (anonymous):

\[log((7-x)/(3x+2))=1\] \[log((7-x)/(3x+2))= log10\] \[(7-x)/(3x+2)= 10\] \[x= -13/31\]